# Scientific Machine Learning of Chaotic Systems Discovers Governing Equations for Neural Populations

### v2

@@ Abstract
Discovering governing equations that describe complex chaotic systems remains a fundamental challenge in physics and neuroscience. Here, we introduce a method, the PEM-UDE, that combines the prediction-error method with universal differential equations to extract interpretable mathematical expressions from chaotic dynamical systems, even with limited or noisy observations. This approach succeeds where traditional techniques fail by smoothing optimization landscapes and removing the chaotic properties during the fitting process without distorting optimal parameters. We demonstrate its efficacy by recovering hidden states in the Rössler system and reconstructing dynamics from noise-corrupted electrical circuit data, where the correct functional form of the dynamics are recovered even when one of the observed time series is corrupted by noise 5x the magnitude of the true signal. We showcase that this method with prior knowledge embeddings is able to recover the correct dynamics where direct symbolic regression methods, such as STLSQ, fail with the given amount of data and noise. Importantly, when applied to neural populations, our method derives novel governing equations that respect biological constraints such as network sparsity – a constraint necessary for cortical information processing yet not captured in next-generation neural mass models – while preserving microscale neuronal parameters. These equations predict an emergent relationship between connection density and both oscillation frequency and synchrony in neural circuits. We validate these predictions using three intracranial electrode recording datasets from the medial entorhinal cortex, prefrontal cortex, and orbitofrontal cortex. Our work provides a pathway to develop mechanistic, multi-scale brain models that generalize across diverse neural architectures, bridging the gap between single-neuron dynamics and macroscale brain activity.
@@ I Introduction
Discovering the equations that govern complex systems has been the goal of dynamics for much of the past century and frequently requires both an expert knowledge of the domain of application and a fair understanding of statistical mechanics to achieve [1, 2, 3, 4]. When successful, this approach has led to significant advancements in our ability to model and predict the behavior of complex systems in applications ranging from climate models [5, 6] to chemical reactions [7, 8] to human social interactions [9]. Recent advances in complex systems theory rely on automated approaches to equation discovery, which have emerged to assist the normally expert-driven process of describing data with systems of differential equations [10, 11, 12, 13]. Symbolic regression techniques, including sparse identification methods (SINDy; [10]) and genetic algorithm approaches [14], have been developed to fit equations directly from data. Although promising, these approaches are limited by their difficulties fitting noisy data [15]. More recently, Universal Differential Equations (UDEs) have emerged as a powerful approach that combines the strengths of machine learning with traditional scientific computing techniques [16]. UDEs incorporate artificial neural networks within differential equation frameworks to learn unknown dynamics directly from data. This approach relies on training a universal approximator under two constraints: imposing a known functional form on the dynamics before learning additional unknown terms [16], and imposing structural constraints on the learning process itself [17].
@@
While useful, these approaches struggle when fitting chaotic systems, which are ubiquitous in neuroscience and other biological contexts [18, 19, 20, 21]. Even simple physical systems can exhibit complex, chaotic solutions when certain parameters are brought past a bifurcation point [22]. Neural dynamics, in particular, often display characteristics of chaotic oscillators [20, 23]. The presence of chaos introduces numerical problems in the fitting, as the trajectory of a chaotic system is not well-defined beyond a Lyapunov time. Formally, any numerical ODE solver introduces truncation error in each step and, due to sensitive dependence on inputs, that ODE solver output will have an $O(1)$ error in the produced trajectory shortly after a Lyapunov time. The only guarantee that the numerical simulation gives a meaningful trajectory is provided by the shadowing lemma, which states that there exists an epsilon ball around the initial condition in which the numerical solver’s trajectory matches the trajectory of one of the other potential initial conditions within the epsilon ball (a statement of density of solutions around the attractor) [24]. The shadowing lemma introduces two crucial factors that make it a particularly challenging problem to learn. The divergent nature of chaotic trajectories creates fundamental instability in all sensitivity analysis methods used to compute gradients. This instability means that standard automatic differentiation through solvers produces inaccurate derivative estimates unless specialized modifications are applied [25, 26, 27, 28, 29]. As a consequence, the standard method for calibration of trajectories, i.e. fitting a loss function defined by numerical solutions of the ODE solver, is not a convergent algorithm in the context of chaotic systems. Additionally, noisy observational data can create a loss function that cannot distinguish between nearby trajectories during training, leading to inaccuracies in the fit dynamics [24]. In other words, with exponential dependence of the solution on initial parameters, infinitesimal changes to the data may land the data point exponentially further along the time series trajectory. Because of these properties, the long-term time series of a chaotic system can only be interpreted probabilistically, based on statistical or ergodic properties [30, 31, 32]. This makes the traditional techniques of fitting a time series via the L2 norm simply ill-defined in these chaotic contexts. Finally, in real-world models, we additionally only have partial observability, which makes parameter fitting for any dynamical system challenging. This is particularly true in neuroscience, where neural interactions form complex systems operating across multiple scales—from molecular processes like neurotransmitter activity to brain-wide electric fields that can be observed on EEG [33, 19, 34].
@@
Aside from the complexity inherent in describing these multi-scale interactions, the dynamics of the brain are often poised near “critical points” that allow for adaptive switching between different dynamic states depending on task requirements, metabolic resources available, and overall health of the neural circuits [35, 36, 37]. While there is an advantage to flexibility and information processing by being located at these critical points, this means that neural dynamics are intrinsically sensitive to small perturbations [38, 20]. Traditional approaches to modeling neuron populations, termed neural mass models, focused on capturing specific macroscale properties without a systematic approach to accounting for microscale biophysical parameters (e.g., ion channel conductance) [39, 40, 41]. More recently, next-generation neural mass models have been shown to provide an analytical expression for population-level activity that preserves the microscale detail present in individual neurons while also capturing emergent macroscale properties (e.g., time-varying synchrony) [42, 43, 44]. These analytical approaches typically rely on the Ott-Antonsen ansatz [45] to avoid a Fokker-Planck derivation of the mean-field dynamics [46, 47]. An important caveat of these derivations is that while they capture many desirable multi-scale dynamics, they rely on non-biophysical assumptions (e.g., all-to-all connectivity between neurons [42], current conservation throughout the system [47]). Although neuron-scale simulations that violate these assumptions are possible, it is difficult or impossible to analytically account for these violations in new derivations of neural masses, thereby limiting their applications to full-scale brain models.
@@
The fundamental disconnect between the underlying neurobiology and the assumption of fully-connected networks that underlies these derivations presents both a challenge and an opportunity. While this assumption is approximately true for deep and subcortical structures [43], it is far from physiological in the cortex, where only 1-5% connectivity is the norm [48]. This is also not an evolutionary afterthought or accident, as there are converging streams of evidence that the sparse, structured connectivity patterns exhibited in the cortex are essential for structured information flow that gives rise to higher cognitive processes [49, 50]. Accurately accounting for sparsity in equations that capture population activity is therefore a necessity in developing accurate models of the brain.
@@
To overcome these limitations, we present a novel method for training UDEs to fit chaotic system dynamics. We apply this method to derive novel next-generation neural mass models that generalize from subcortical to cortical dynamics. We begin by describing the procedure, which leverages the prediction-error method (PEM) [17, 51] to assist in training the UDEs on chaotic system data, and then derives a symbolic form of the learned expression from the UDE to generalize the dynamics. We demonstrate the ability of this approach to learn unknown states in a chaotic system, showing that it can accurately recover a hidden state in the Rössler system that traditional UDE approaches cannot and that it does so by removing the sensitive dependence on parameter choices from the system. We then demonstrate that our approach learns the dynamics of a chaotic electrical circuit even under incomplete observation of states, where a dimension of the observed data is corrupted by noise 5x the original signal. Notably, we showcase that the PEM-UDE approach is more data-efficient than pure symbolic regression by demonstrating that sequential thresholding least squares (STLSQ) is unable to recover the correct dynamic system under these conditions. Finally, we demonstrate the system’s ability to learn the dynamics of a neural mass model directly from the firing rate and voltage of a network of spiking neurons, showing that it can successfully learn the mean-field dynamics from data that are typically the only states observable experimentally. Our approach demonstrates that the symbolic form we have developed generalizes across different brain regions more effectively than traditional neural mass models. Our modeling directly reveals how varying connectivity structures throughout the brain produce distinct rhythmic patterns in different regions.
@@ II.1.1 Overview
A graphical overview of the approach is provided in Fig. 1. We begin with observations of a physical or biological system that result in a dataset of related time series. Ideally, we would like to have a complete system of differential equations that describes the dynamics of the system under observation and makes accurate predictions of future dynamics. Based upon these observations and previous work on similar systems, we can create an initial guess of a description of the dynamics, leveraging prior information from physical systems to structure the first approximation of the system’s behavior [16]. We then set up a UDE system to learn the unknown dynamics directly from the observed data. Traditional methods for training UDEs struggle to accurately fit chaotic systems, as shown in the results. To address this limitation, we developed a novel training approach that integrates PEM dynamics with the UDE during simulations, enabling accurate learning from the data. Once training is complete, we extract a symbolic representation of the dynamics from the UDE network using either sparse linear solver based symbolic regression techniques (STLSQ) or with evolutionary algorithm based symbolic regression methods [14], depending on the specific context and data availability. This symbolic description allows us to both predict future time series and make analytical observations about the underlying dynamics.
@@ II.1.2 UDE approach
Following prior work [16], a universal differential equation (UDE) with an additive correction of a general process $\mathbf{x}(t)$ with some observable state $\mathbf{y}(t)$ is given by
@@
Here we use bold variables to denote vectors. $f_{p}$ are the known (or assumed) dynamics of the system with parameters $p$, $\mathbf{x}_{p,\theta}(t;\mathbf{x}_{0})$ denotes the solution of those dynamics with initial conditions $\mathbf{x}_{0}$ at parameter values $p$ and $\theta$, $\mathbf{u}(t)$ is an external stimulus, $h_{q}(\mathbf{x}_{p,\theta}(t;\mathbf{x}_{0}),t)$ is the observation function parameterized by $q$ which can explicitly depend on time (for instance, due to some noise processes or the influence of the external stimulus on the observation etc.), and $NN_{\theta}(\mathbf{x})$ is the universal approximator network trained to learn the unknown dynamics of the system based on its parameters $\theta$. Lastly, note that for notational simplicity we drop the parameter dependence of the observation $\mathbf{y}_{\theta}(t)$ on parameters $p$ and $q$ and initial conditions $\mathbf{x}_{0}$, but keep the network parameters $\theta$ since we want to optimize those while keeping the rest fixed. While this work focuses on systems of ordinary differential equations, it is not limited to these scenarios, and can be extended to stochastic and delayed differential equation systems as well, as has been done in previous work using UDEs [16]. For some observed dataset $\{\hat{\mathbf{y}}_{i}\}_{i=1,\dots,N}$, training the UDE $NN_{\theta}$ will use a cost function $C(\theta)$ that minimizes the error at every observed time point $t_{i}$ such that
@@ II.1.3 PEM approach
The prediction-error method (PEM) is a standard optimization technique in fitting parameters of dynamical systems to observational data [17, 52]. Similarly to Eq. (1), we consider a general process $\mathbf{x}(t)$ that gives rise to some observable state $\mathbf{y}(t)$ with some measurement noise $\mathbf{e}(t)$
@@
In this statement, $\mathbf{x}$ is the state of a time-varying system with parameters $p$ that we wish to fit. The observed state $\mathbf{y}$ is defined by a measurement function $m$ of the process $\mathbf{x}(t)$, parameterized by $q$, and a measurement error $\mathbf{e}(t)$. Here, we assume that there is a one-to-one mapping of $\mathbf{x}$ to $\mathbf{y}$ (i.e., there are no terms in $\mathbf{y}$ that depend on multiple elements of $\mathbf{x}$). For a more general discussion of the conditions under which PEM can be applied, see [17, 52]. For example, $m_{q,i}$ is not limited to being a linear function of $x_{i}$, but can also perform a convolution or incorporate delays [52]. In all of our examples, a few components of $m_{q,i}$ are zero, indicating that these components, $y_{p,i}$, are not observed. The estimation error $\epsilon$ is weighted by a function $K(t,\beta)$ and is added at each time step to the original dynamics $\mathbf{x}(t)$ to create an adjusted estimate $\tilde{\mathbf{x}}(t)$ and observations $\tilde{\mathbf{y}}(t)$ such that
@@
where, as defined above, $\hat{\mathbf{y}}(t)$ is the observed data, interpolated to whatever step is needed by the solver. This error gain factor $K(t,\beta)$ can take on different forms. In this work, we set $K(t,\beta)=K$, thereby employing the simplest possible form of what previous work has called the parameterized observer [52]. If there is strong reason to suspect that the quality of the observations varies over time (e.g. significantly increased noise present at certain time intervals), then including this information in a time-varying $K(t,\beta)$ for instance by expanding the equations (4) to resemble an extended Kalman filter (where $K(t,\beta)$ is the Kalman gain), could provide better performance when fitting. However, we note that previous work has shown that the parameterized observer outperforms the extended Kalman filter in practice, as tuning the hyperparameter over time is difficult [52]. Although for notational simplification we will use $K$ in the remaining sections, this is done without loss of generality, as implementing an extended Kalman filter can easily be done if there is a data-driven reason to require it.

### v3

@@ Abstract
Extracting interpretable mathematical models from chaotic dynamical systems represents a fundamental challenge in physics and neuroscience, where traditional equation discovery methods fail due to sensitive dependence on initial conditions and observational noise. We introduce PEM-UDE, a method that combines the prediction-error method with universal differential equations to discover governing equations from chaotic time series data. The key innovation lies in using prediction-error feedback to smooth otherwise intractable chaotic optimization landscapes during neural network training, effectively removing sensitive dependence on parameters while preserving the location of the global minimum. We demonstrate the method’s capability by recovering hidden dynamics in the Rössler system for which standard UDE approaches fail, and by reconstructing correct functional forms from experimental electrical circuit data corrupted by noise five times the signal magnitude. Following UDE training, we extract symbolic expressions using sparse regression techniques, enabling analytical interpretation and generalization beyond the training data. Importantly, this two-stage approach succeeds where direct symbolic regression methods fail on the same noisy chaotic datasets. Applied to neuroscience, PEM-UDE yields novel next-generation neural mass models that incorporate biologically realistic network sparsity - a critical constraint for cortical information processing, which is absent from existing analytical derivations that assume all-to-all connectivity. The discovered equations predict emergent relationships between synaptic connection density, oscillation frequency, and network synchronization, which we validate using three independent intracranial electrode datasets from the rat medial entorhinal cortex, rat prefrontal cortex, and human orbitofrontal cortex. These predictions accurately capture the transition from theta-dominant rhythms in densely connected regions to alpha-dominant rhythms in sparse cortical areas, along with corresponding changes in baseline synchrony levels. Our framework provides a general solution for mechanistic model discovery in complex systems, where chaos and noise render traditional approaches ineffective. It offers immediate applications across physics, engineering, and computational neuroscience, bridging microscale dynamics with macroscale emergent phenomena.
@@ I Introduction
Discovering the equations that govern complex systems has been the goal of dynamics for much of the past century and frequently requires both an expert knowledge of the domain of application and a fair understanding of statistical mechanics to achieve [schmidt2009, king2009, cranmer2020, feinerman2018]. When successful, this approach has led to significant advancements in our ability to model and predict the behavior of complex systems in applications ranging from climate models [kochkov2024, foster2020] to chemical reactions [kim2021, ziepke2022] to human social interactions [pagan2019]. Recent advances in complex systems theory rely on automated approaches to equation discovery, which have emerged to assist the normally expert-driven process of describing data with systems of differential equations [brunton2016, beregi2023, udrescu2020, ellis2023]. Symbolic regression techniques, including sparse identification methods (SINDy; [brunton2016]) and genetic algorithm approaches [cranmer2023], have been developed to fit equations directly from data. Although promising, these approaches are limited by their difficulties fitting noisy data [bakarji2023a]. More recently, Universal Differential Equations (UDEs) have emerged as a powerful approach that combines the strengths of machine learning with traditional scientific computing techniques [rackauckas2021]. UDEs incorporate artificial neural networks within differential equation frameworks to learn unknown dynamics directly from data. This approach relies on training a universal approximator under two constraints: imposing a known functional form on the dynamics before learning additional unknown terms [rackauckas2021], and imposing structural constraints on the learning process itself [ljung2002].
@@
While useful, these approaches struggle when fitting chaotic systems, which are ubiquitous in neuroscience and other biological contexts [hannay2018, breakspear2017, chesebro2023, abrevaya2024effective]. Even simple physical systems can exhibit complex, chaotic solutions when certain parameters are brought past a bifurcation point [rossler1976]. Neural dynamics, in particular, often display characteristics of chaotic oscillators [chesebro2023, schmidt2020]. The presence of chaos introduces numerical problems in the fitting, as the trajectory of a chaotic system is not well-defined beyond a Lyapunov time. Formally, any numerical ODE solver introduces truncation error in each step and, due to sensitive dependence on inputs, that ODE solver output will have an $O(1)$ error in the produced trajectory shortly after a Lyapunov time. The only guarantee that the numerical simulation gives a meaningful trajectory is provided by the shadowing lemma, which states that there exists an epsilon ball around the initial condition in which the numerical solver’s trajectory matches the trajectory of one of the other potential initial conditions within the epsilon ball (a statement of density of solutions around the attractor) [chandramoorthy2021]. The shadowing lemma introduces two crucial factors that make it a particularly challenging problem to learn. The divergent nature of chaotic trajectories creates fundamental instability in all sensitivity analysis methods used to compute gradients. This instability means that standard automatic differentiation through solvers produces inaccurate derivative estimates unless specialized modifications are applied [errico_what_1997, ni_adjoint_2019, ni_sensitivity_2019, chater_least_2017, wang_least_2014]. As a consequence, the standard method for calibration of trajectories, i.e. fitting a loss function defined by numerical solutions of the ODE solver, is not a convergent algorithm in the context of chaotic systems. Additionally, noisy observational data can create a loss function that cannot distinguish between nearby trajectories during training, leading to inaccuracies in the fit dynamics [chandramoorthy2021]. In other words, with exponential dependence of the solution on initial parameters, infinitesimal changes to the data may land the data point exponentially further along the time series trajectory. Because of these properties, the long-term time series of a chaotic system can only be interpreted probabilistically, based on statistical or ergodic properties [eckmann_ergodic_1985, petersen_ergodic_1989, bradley_nonlinear_2015]. This makes the traditional techniques of fitting a time series via the L2 norm simply ill-defined in these chaotic contexts. Finally, in real-world models, we additionally only have partial observability, which makes parameter fitting for any dynamical system challenging. This is particularly true in neuroscience, where neural interactions form complex systems operating across multiple scales—from molecular processes like neurotransmitter activity to brain-wide electric fields that can be observed on EEG [adam2024, breakspear2017, antal2024].
@@
Aside from the complexity inherent in describing these multi-scale interactions, the dynamics of the brain are often poised near “critical points” that allow for adaptive switching between different dynamic states depending on task requirements, metabolic resources available, and overall health of the neural circuits [weistuch2021, habibollahi2023, tian2022]. While there is an advantage to flexibility and information processing by being located at these critical points, this means that neural dynamics are intrinsically sensitive to small perturbations [graf2024, chesebro2023]. Traditional approaches to modeling neuron populations, termed neural mass models, focused on capturing specific macroscale properties without a systematic approach to accounting for microscale biophysical parameters (e.g., ion channel conductance) [wilson1972, jansen1995, larter1999]. More recently, next-generation neural mass models have been shown to provide an analytical expression for population-level activity that preserves the microscale detail present in individual neurons while also capturing emergent macroscale properties (e.g., time-varying synchrony) [montbrio2015, chen2022, coombes2019]. These analytical approaches typically rely on the Ott-Antonsen ansatz [ott2008a] to avoid a Fokker-Planck derivation of the mean-field dynamics [goldobin2024, divolo2018]. An important caveat of these derivations is that while they capture many desirable multi-scale dynamics, they rely on non-biophysical assumptions (e.g., all-to-all connectivity between neurons [montbrio2015], current conservation throughout the system [divolo2018]). Recent work has extended these approaches to account for fluctuations arising from sparse connectivity and noise in networks of quadratic integrate-and-fire neurons [goldobin2021], representing important progress. However, such extensions remain limited to simplified neuron models lacking biophysically-detailed adaptation and recovery dynamics. More broadly, extending analytical mean-field derivations to biophysically detailed neuron models (e.g., Izhikevich neurons with adaptation currents) while simultaneously relaxing connectivity assumptions remains an open challenge, thereby motivating data-driven approaches.
@@
The fundamental disconnect between the underlying neurobiology and the assumption of fully-connected networks that underlies these derivations presents both a challenge and an opportunity. While this assumption is approximately true for deep and subcortical structures [chen2022], it is far from physiological in the cortex, where only 1-5% connectivity is the norm [pathak2024]. This is also not an evolutionary afterthought or accident, as there are converging streams of evidence that the sparse, structured connectivity patterns exhibited in the cortex are essential for structured information flow that gives rise to higher cognitive processes [balcioglu2023, fruengel2025]. Accurately accounting for sparsity in equations that capture population activity is therefore a necessity in developing accurate models of the brain.
@@
To overcome these limitations, we present a novel method for training UDEs to fit chaotic system dynamics. We apply this method to derive novel next-generation neural mass models that generalize from subcortical to cortical dynamics. We begin by describing the procedure, which leverages the prediction-error method (PEM) [ljung2002, stoica1989] to assist in training the UDEs on chaotic system data, and then derives a symbolic form of the learned expression from the UDE to generalize the dynamics. We demonstrate the ability of this approach to learn unknown states in a chaotic system, showing that it can accurately recover a hidden state in the Rössler system that traditional UDE approaches cannot and that it does so by removing the sensitive dependence on parameter choices from the system. We then demonstrate that our approach learns the dynamics of a chaotic electrical circuit even under incomplete observation of states, where a dimension of the observed data is corrupted by noise 5x the original signal. Notably, we showcase that the PEM-UDE approach is more data-efficient than pure symbolic regression by demonstrating that sequential thresholding least squares (STLSQ) is unable to recover the correct dynamic system under these conditions. Finally, we demonstrate the system’s ability to learn the dynamics of a neural mass model directly from the firing rate and voltage of a network of spiking neurons, showing that it can successfully learn the mean-field dynamics from data that are typically the only states observable experimentally. Our approach demonstrates that the symbolic form we have developed generalizes across different brain regions more effectively than traditional neural mass models. Our modeling directly reveals how varying connectivity structures throughout the brain produce distinct rhythmic patterns in different regions.
@@ II.1.1 Overview
A graphical overview of the approach is provided in Fig. 1. We begin with observations of a physical or biological system that result in a dataset of related time series. Ideally, we would like to have a complete system of differential equations that describes the dynamics of the system under observation and makes accurate predictions of future dynamics. Based upon these observations and previous work on similar systems, we can create an initial guess of a description of the dynamics, leveraging prior information from physical systems to structure the first approximation of the system’s behavior [rackauckas2021]. We then set up a UDE system to learn the unknown dynamics directly from the observed data. Traditional methods for training UDEs struggle to accurately fit chaotic systems, as shown in the results. To address this limitation, we developed a novel training approach that integrates PEM dynamics with the UDE during simulations, enabling accurate learning from the data. Once training is complete, we extract a symbolic representation of the dynamics from the UDE network using either sparse linear solver-based symbolic regression techniques (STLSQ) or with evolutionary algorithm-based symbolic regression methods [cranmer2023], depending on the specific context and data availability. This symbolic description allows us to both predict future time series and make analytical observations about the underlying dynamics.
@@ II.1.2 UDE approach
Following prior work [rackauckas2021], a universal differential equation (UDE) with an additive correction of a general process $\mathbf{x}(t)$ with some observable state $\mathbf{y}(t)$ is given by
@@
Here we use bold variables to denote vectors. $f_{p}$ are the known (or assumed) dynamics of the system with parameters $p$, $\mathbf{x}_{p,\theta}(t;\mathbf{x}_{0})$ denotes the solution of those dynamics with initial conditions $\mathbf{x}_{0}$ at parameter values $p$ and $\theta$, $\mathbf{u}(t)$ is an external stimulus, $h_{q}(\mathbf{x}_{p,\theta}(t;\mathbf{x}_{0}),t)$ is the observation function parameterized by $q$ which can explicitly depend on time (for instance, due to some noise processes or the influence of the external stimulus on the observation etc.), and $NN_{\theta}(\mathbf{x})$ is the universal approximator network trained to learn the unknown dynamics of the system based on its parameters $\theta$. Lastly, note that for notational simplicity we drop the parameter dependence of the observation $\mathbf{y}_{\theta}(t)$ on parameters $p$ and $q$ and initial conditions $\mathbf{x}_{0}$, but keep the network parameters $\theta$ since we want to optimize those while keeping the rest fixed. While this work focuses on systems of ordinary differential equations, it is not limited to these scenarios, and can be extended to stochastic and delayed differential equation systems as well, as has been done in previous work using UDEs [rackauckas2021]. For some observed dataset $\{\hat{\mathbf{y}}_{i}\}_{i=1,\dots,N}$, training the UDE $NN_{\theta}$ will use a cost function $C(\theta)$ that minimizes the error at every observed time point $t_{i}$ such that
@@ II.1.3 PEM approach
The prediction-error method (PEM) is a standard optimization technique in fitting parameters of dynamical systems to observational data [ljung2002, larsson2009]. Similarly to Eq. (1), we consider a general process $\mathbf{x}(t)$ that gives rise to some observable state $\mathbf{y}(t)$ with some measurement noise $\mathbf{e}(t)$
@@
In this statement, $\mathbf{x}$ is the state of a time-varying system with parameters $p$ that we wish to fit. The observed state $\mathbf{y}$ is defined by a measurement function $m$ of the process $\mathbf{x}(t)$, parameterized by $q$, and a measurement error $\mathbf{e}(t)$. Here, we assume that there is a one-to-one mapping of $\mathbf{x}$ to $\mathbf{y}$ (i.e., there are no terms in $\mathbf{y}$ that depend on multiple elements of $\mathbf{x}$). For a more general discussion of the conditions under which PEM can be applied, see [ljung2002, larsson2009]. For example, $m_{q,i}$ is not limited to being a linear function of $x_{i}$, but can also perform a convolution or incorporate delays [larsson2009]. In all of our examples, a few components of $m_{q,i}$ are zero, indicating that these components, $y_{p,i}$, are not observed. The estimation error $\epsilon$ is weighted by a function $K(t,\beta)$ and is added at each time step to the original dynamics $\mathbf{x}(t)$ to create an adjusted estimate $\tilde{\mathbf{x}}(t)$ and observations $\tilde{\mathbf{y}}(t)$ such that
@@
where, as defined above, $\hat{\mathbf{y}}(t)$ is the observed data, interpolated to whatever step is needed by the solver. This error gain factor $K(t,\beta)$ can take on different forms. In this work, we set $K(t,\beta)=K$, thereby employing the simplest possible form of what previous work has called the parameterized observer [larsson2009]. If there is strong reason to suspect that the quality of the observations varies over time (e.g. significantly increased noise present at certain time intervals), then including this information in a time-varying $K(t,\beta)$ for instance by expanding the equations (4) to resemble an extended Kalman filter (where $K(t,\beta)$ is the Kalman gain), could provide better performance when fitting. However, we note that previous work has shown that the parameterized observer outperforms the extended Kalman filter in practice, as tuning the hyperparameter over time is difficult [larsson2009]. Although for notational simplification we will use $K$ in the remaining sections, this is done without loss of generality, as implementing an extended Kalman filter can easily be done if there is a data-driven reason to require it.